Question text


Madison usually receives 4 letters in the mail each week. Find the probability of receiving 8 letters for the next 2 weeks.


1
Expert's answer
2022-03-10T16:21:59-0500

This is a Poisson distribution with "\mu=4*2=8."

"P(X=8)=e^{-8}\frac{8^8}{8!}=0.1396".